OpenStudy (anonymous):

determine whether the ordered pair is a solution of the inequality: (-3,4); 9x+y<-8 In this case is the ordered pair a solution? yes or no

OpenStudy (anonymous):

determine whether the ordered pair is a solution of the inequality: (-3,4); 9x+y<-8 In this case is the ordered pair a solution? yes or no

OpenStudy (anonymous):

substitute values. If LS<RS then it is a solution -27+4 is less than -8

OpenStudy (anonymous):

can you show the work

OpenStudy (anonymous):

work is exactly what mathguy wrote. replace x by -3, y by 4 and see if your result it less than -8. you write \[9\times -3 + 4<-8\] \[-27+4,-8\] \[-23< -8\] true!

OpenStudy (anonymous):

thanks you guys

OpenStudy (anonymous):

7x + 8 > or = to - 48 or 5x + 2 > or = to - 8 the solution of the compound inequality is {x|x>=to ---?